# Information-Theoretic Constraints on Variational Quantum Optimization: Efficiency Transitions and the Dynamical Lie Algebra

### v1

@@ Abstract
Variational quantum algorithms are the leading candidates for near-term quantum advantage, yet their scalability is limited by the “Barren Plateau” phenomenon. While traditionally attributed to geometric vanishing gradients, we propose an information-theoretic perspective. Using ancilla-mediated coherent feedback, we demonstrate an empirical constitutive relation $\Delta E\leq\eta I(S:A)$ linking work extraction to mutual information, with quantum entanglement providing a factor-of-2 advantage over classical Landauer bounds. By scaling the system size, we identify a distinct efficiency transition governed by the dimension of the Dynamical Lie Algebra. Systems with polynomial algebraic complexity exhibit sustained positive efficiency, whereas systems with exponential complexity undergo an “efficiency collapse” ($\eta\to 0$) at $N\approx 6$ qubits. These results suggest that the trainability boundary in variational algorithms correlates with information-theoretic limits of quantum feedback control.
@@ I Introduction
Recent theoretical advances have identified the Dynamical Lie Algebra (DLA) as the key structural predictor of trainability [ragone2024lie]. Specifically, circuits generating a polynomially scaling DLA are proven to escape barren plateaus, while those generating exponentially scaling DLAs succumb to them. We leverage this algebraic classification to probe the thermodynamic stability of these distinct regimes.
Here, we propose that computational hardness in quantum circuits correlates with information-theoretic constraints on feedback control. By treating the ancilla interaction as an effective environment, we recover non-linear control dynamics via information erasure, consistent with open-system decoherence approaches to measurement [zurek2003decoherence].
We reframe the variational optimizer not as a mathematical function, but as a quantum Maxwell’s Demon. To isolate the thermodynamic contribution, we implement a decoupled ’Coherent Feedback’ protocol that fixes the actuation strength while varying the sensing duration, thereby proving that the optimization work is causally driven by the mutual information channel capacity. In this framework, the feedback loop acts as a thermodynamic engine: it extracts entropy from the system to lower its Hamiltonian expectation value $\langle H\rangle$ (Work), fueled by the mutual information established between the system and the control ancilla.
@@
Here, we extend this principle to the thermodynamic limit of many-body optimization. We ask: Does the capacity to extract work scale indefinitely, or does it face a critical complexity barrier? By comparing systems with polynomial vs. exponential algebraic complexity, we demonstrate that information-driven work extraction is not a guaranteed resource but is strictly bounded by the algebraic structure of the problem Hamiltonian.
Crucially, this ancilla-mediated feedback induces an effective non-linearity in the parameter update trajectory. We demonstrate this via a controlled experiment comparing interacting vs. non-interacting Hamiltonians (see Methods). For a separable Hamiltonian $H=Z_{0}+Z_{1}$, where the phase factors as $e^{-i(E_{0}+E_{1})t}=e^{-iE_{0}t}\cdot e^{-iE_{1}t}$, we observe zero mutual information between parameter qubits:
@@
In contrast, for an interacting Hamiltonian $H=Z_{0}Z_{1}$, where the phase depends on the product $E_{0}\cdot E_{1}$, we observe significant entanglement:
@@
By analyzing the scaling of this feedback mechanism, we identify a distinct efficiency transition governed by the dimension of the Dynamical Lie Algebra (DLA). We observe that systems with polynomial algebraic complexity (Ordered phase) exhibit “constructive information scaling,” where efficiency increases with system size. In contrast, systems with exponential complexity (Chaotic phase) undergo an “Efficiency Collapse.” This suggests that the trainability boundary in variational algorithms correlates with the information channel capacity of the feedback controller.
Alternative strategies, such as the Quantum Walk-based Optimization Algorithm (QWOA) [glos2019quantum, marsh2020quantum], have sought to overcome local minima by exploiting coherent tunneling and variable-time evolution. Our framework unifies the QWOA paradigm with variational optimization by showing that the ancilla acts as a quantum walk “coin qubit” controlling trap-diffusion dynamics. While these dynamic approaches improve exploration, the fundamental thermodynamic bounds governing their efficiency in the many-body limit—specifically the transition from coherent flow to information scrambling—remain uncharacterized.
@@ II Thermodynamic Constitutive Relations
We first establish the source of non-linearity via a minimal 4-qubit experiment. For a non-interacting Hamiltonian $H=Z_{0}+Z_{1}$, the evolution phase factors separably, yielding zero correlation between parameter qubits (Eq. 1). For an interacting Hamiltonian $H=Z_{0}Z_{1}$, the coupled phase creates entanglement, yielding $I(P_{0}:P_{1})=1.74$ bits (Eq. 2). This demonstrates that non-linearity requires interaction terms that couple parameter subspaces—a necessary condition for the trap-diffusion mechanism to operate.
@@ II.1 The Trap-Diffusion Mechanism
The core insight connecting our protocol to quantum walk speedup lies in the controlled-operation structure. In our W-gate protocol (see Methods), the circuit applies:
@@
This mechanism is mathematically equivalent to Grover’s amplitude amplification [grover1996fast], where the oracle marks “good” states (low energy) and the diffuser redistributes amplitude. Our protocol realizes this as a continuous process: the controlled-mixer acts as the diffuser, and the energy-dependent sensing acts as a soft oracle that preferentially marks high-energy states for trapping.
Crucially, we observe that the work extraction correlates with Logarithmic Negativity ($R\approx 0.9$), a strict entanglement monotone. A purely classical feedback loop (zero entanglement) would yield zero work in this protocol, confirming the quantum nature of the mechanism.
To rigorously characterize the thermodynamic cost of this non-linearity, we implemented a controlled “Coherent Feedback” protocol (see Methods) designed to decouple the information gathering phase from the feedback actuation. Unlike standard variational updates where parameters are modified classically, our protocol utilizes a coherent conditional rotation ($CR_{X}$) acting directly on the system, triggered by the ancilla’s state.
Figure 1: The Thermodynamic Constitutive Law. Extracted Work ($\Delta\langle H\rangle$) vs. Mutual Information $I(S:A)$ for a 4-qubit transverse Ising system. The robust linear relationship ($R^{2}\approx 0.90$, slope $\eta=0.247$ energy/bit) confirms the constitutive law $\Delta E\leq\eta I$. The sensing time $\tau$ was varied from 0 to 1.5 while the feedback strength was held constant at $\theta_{gain}=0.5$ rad, isolating information as the sole variable driving work extraction.
Crucially, we fixed the feedback actuation strength (the “kick”) to a constant value and varied only the sensing duration $dt$. This isolates the information content as the sole variable driving the performance difference. We observe that the extracted work ($\Delta\langle H\rangle$) scales linearly with the mutual information $I(S:A)$ established between the system and the ancilla, yielding a strict constitutive Equation of State:
@@
Here, the proportionality constant $\eta$ represents the “Algorithmic Efficiency” or informational conductance of the ansatz-Hamiltonian pair. Our data indicates a robust linear correlation ($R^{2}\approx 0.90$), confirming that the optimization rate is fundamentally limited by the information channel capacity of the probe, rather than the driving force of the optimizer.
To determine the physical nature of these correlations, we measured the Logarithmic Negativity $E_{N}$, a strict monotone of quantum entanglement which is zero for all separable (classical) states. We find that the work extraction correlates directly with the generated entanglement:
@@
The persistence of this correlation ($R\approx 0.90$) and the strictly non-zero value of $E_{N}$ confirms that the mechanism relies specifically on quantum interference effects. This distinguishes the VQA optimizer from a classical Szilard engine, establishing that quantum entanglement is the essential thermodynamic resource driving the cooling process.
To verify that this mechanism provides genuine thermodynamic advantage, we computed the Landauer erasure cost. Remarkably, we observe a constant ratio $I(S:A)/S(A)=2.00$ across all sensing times (Fig. 2). This ratio of exactly 2 is the hallmark of pure bipartite entanglement, where $I(S:A)=2S(A)$. Since the Landauer cost for erasing the ancilla is $W_{cost}=k_{B}T\cdot S(A)$, while the extractable work scales as $W_{ext}\propto I(S:A)$, the quantum correlations provide a factor-of-2 advantage over the classical limit, yielding net positive work after accounting for erasure.
Figure 2: Quantum Advantage via Landauer Analysis. Extracted work (blue) vs. Landauer erasure cost (red dashed) as a function of sensing time $\tau$. The green shaded region indicates net positive work after accounting for information erasure. The constant ratio $I(S:A)/S(A)=2.00$ confirms that quantum entanglement (not classical correlation) fuels the engine.
@@ III The Complexity-Dependent Efficiency Transition
To investigate the limits of this feedback mechanism, we analyzed the scaling of the algorithmic efficiency $\eta$ across two distinct topological classes of Hamiltonians as a function of system size $N$. We compared an “Ordered” system (Complete Graph $K_{n}$) characterized by a polynomial Dynamical Lie Algebra (DLA) dimension ($O(N^{3})$) [allcock2024dynamical] against a “Chaotic” system (Sherrington-Kirkpatrick Spin Glass) characterized by an exponential DLA dimension ($O(4^{N})$) [larocca2023diagnosing]. While the system sizes ($N\leq 8$) are limited by the exponential cost of classical simulation, the observed efficiency collapse in chaotic systems is consistent with the ancilla channel capacity becoming insufficient to resolve the scrambled gradient information.
The Lie algebraic classification of trainability was recently unified by Ragone et al. [ragone2024lie], who proved the exact variance formula $\text{Var}[\ell]=\mathcal{P}_{\mathfrak{g}}(\rho)\cdot\mathcal{P}_{\mathfrak{g}}(O)/\dim(\mathfrak{g})$, establishing that $\dim(\mathfrak{g})\in\Omega(b^{n})$ with $b>2$ implies a barren plateau. Our efficiency metric $\eta$ provides a complementary thermodynamic diagnostic: while Ragone et al. predict trainability from gradient variance, we directly measure the information-to-work conversion capacity of the feedback loop. Both observables collapse when DLA dimension grows exponentially, suggesting a unified information-theoretic origin for barren plateaus.
We observe a sharp bifurcation in thermodynamic behavior (Fig. 3). For the Ordered system, we observe a regime of “Constructive Information Scaling,” where the efficiency $\eta$ increases with system size, indicating that the computational resistance to information flow decreases as the Hilbert space grows. In this regime, the growing algebraic structure provides additional pathways for the Demon to navigate the Hilbert space without exceeding its information channel capacity.
Figure 3: The Complexity-Dependent Efficiency Transition. Normalized algorithmic efficiency ($\eta/N^{2}$) vs. System Size NN for Ordered (Ferromagnet, blue) and Chaotic (Spin Glass, red) Hamiltonians. The Ordered phase maintains positive efficiency, while the Chaotic phase undergoes efficiency collapse, with efficiency approaching zero at $N\geq 6$. Error bars represent $\pm 1\sigma$ over 5 random seeds. The transition is consistent with the ancilla channel capacity (1 bit) becoming insufficient to resolve the scrambled gradient information.
@@
In stark contrast, the Chaotic system undergoes an “Efficiency Collapse.” As the system size increases, the efficiency drops precipitously, crossing zero at a critical size $N_{c}\approx 6.4$. At this point, the rate of operator spreading (scrambling) generates new independent gradient directions faster than the single-qubit ancilla (1 bit/cycle) can interrogate them. Beyond $N_{c}$, the information generation rate exceeds the channel capacity, and the ancilla cannot gather sufficient information to guide the trajectory.
To quantify this transition, we introduce the Complexity Specific Heat, defined as the susceptibility of algorithmic efficiency to system size increase:
@@
This quantity is analogous to thermal specific heat $C=\partial\langle E\rangle/\partial T$, measuring the “response” of the optimization engine to changes in problem complexity. In the Ordered phase, $\chi_{\text{comp}}>0$ (efficiency increases with $N$), while in the Chaotic phase, $\chi_{\text{comp}}<0$ and diverges at $N_{c}$, signaling a thermodynamic instability.
This efficiency collapse provides a complementary diagnostic to the Quantum Fisher Information approach of Abbas et al. [abbas2021effective], who showed that the effective dimension $d_{\text{eff}}$ (derived from QFIM eigenvalues) correlates with model trainability. While their metric captures parameter redundancy, our efficiency $\eta$ directly measures the information-to-work conversion capacity of the feedback loop.
@@
For the Ordered phase, we utilized the Complete Graph Hamiltonian ($K_{n}$). Recent algebraic analysis by Allcock et al. [allcock2024dynamical] proves the exact dimension $\dim(\mathfrak{g}_{K_{n}})=\frac{1}{12}(n^{3}+6n^{2}+O(n))$, placing it firmly within the polynomially tractable regime.

### v2

@@ Abstract
Variational quantum algorithms are the leading candidates for near-term quantum advantage, yet their scalability is limited by the “Barren Plateau” phenomenon. While traditionally attributed to geometric vanishing gradients, I propose an information-theoretic perspective. Using ancilla-mediated coherent feedback, I demonstrate an empirical constitutive relation $\Delta E\leq\eta I(S:A)$ linking work extraction to mutual information. The protocol maintains pure bipartite entanglement ($I/S=2$), confirming operation in the quantum correlation regime. By scaling the system size, I identify a distinct efficiency transition governed by the dimension of the Dynamical Lie Algebra. Systems with polynomial algebraic complexity exhibit sustained positive efficiency, whereas systems with exponential complexity undergo an “efficiency collapse” ($\eta\to 0$) at $N\approx 6$ qubits. These results suggest that the trainability boundary in variational algorithms correlates with information-theoretic limits of quantum feedback control.
@@ I Introduction
Recent theoretical advances have identified the Dynamical Lie Algebra (DLA) as the key structural predictor of trainability [ragone2024lie]. Specifically, circuits generating a polynomially scaling DLA are proven to escape barren plateaus, while those generating exponentially scaling DLAs succumb to them. I leverage this algebraic classification to probe the thermodynamic stability of these distinct regimes.
Here, I propose that computational hardness in quantum circuits correlates with information-theoretic constraints on feedback control. By treating the ancilla interaction as an effective environment, I recover non-linear control dynamics via information erasure, consistent with open-system decoherence approaches to measurement [zurek2003decoherence].
I reframe the variational optimizer not as a mathematical function, but as a quantum Maxwell’s Demon. To isolate the thermodynamic contribution, I implement a decoupled ’Coherent Feedback’ protocol that fixes the actuation strength while varying the sensing duration, thereby proving that the optimization work is causally driven by the mutual information channel capacity. In this framework, the feedback loop acts as a thermodynamic engine: it extracts entropy from the system to lower its Hamiltonian expectation value $\langle H\rangle$ (Work), fueled by the mutual information established between the system and the control ancilla.
@@
Here, I extend this principle to the thermodynamic limit of many-body optimization. I investigate whether the capacity for work extraction remains extensive in the large-system limit, or if it undergoes a critical collapse governed by algebraic complexity. By comparing systems with polynomial vs. exponential algebraic complexity, I demonstrate that information-driven work extraction is not a guaranteed resource but is strictly bounded by the algebraic structure of the problem Hamiltonian.
Crucially, this ancilla-mediated feedback induces an effective non-linearity in the parameter update trajectory. I demonstrate this via a controlled experiment comparing interacting vs. non-interacting Hamiltonians (see Methods). For a separable Hamiltonian $H=Z_{0}+Z_{1}$, where the phase factors as $e^{-i(E_{0}+E_{1})t}=e^{-iE_{0}t}\cdot e^{-iE_{1}t}$, I observe zero mutual information between parameter qubits:
@@
In contrast, for an interacting Hamiltonian $H=Z_{0}Z_{1}$, where the phase depends on the product $E_{0}\cdot E_{1}$, I observe significant entanglement:
@@
By analyzing the scaling of this feedback mechanism, I identify a distinct efficiency transition governed by the dimension of the Dynamical Lie Algebra (DLA). I observe that systems with polynomial algebraic complexity (Ordered phase) exhibit “constructive information scaling,” where efficiency increases with system size. In contrast, systems with exponential complexity (Chaotic phase) undergo an “Efficiency Collapse.” This suggests that the trainability boundary in variational algorithms correlates with the information channel capacity of the feedback controller.
Alternative strategies, such as the Quantum Walk-based Optimization Algorithm (QWOA) [glos2019quantum, marsh2020quantum], have sought to overcome local minima by exploiting coherent tunneling and variable-time evolution. My framework unifies the QWOA paradigm with variational optimization by showing that the ancilla acts as a quantum walk “coin qubit” controlling trap-diffusion dynamics. While these dynamic approaches improve exploration, the fundamental thermodynamic bounds governing their efficiency in the many-body limit—specifically the transition from coherent flow to information scrambling—remain uncharacterized.
@@ II Thermodynamic Constitutive Relations
I first establish the source of non-linearity via a minimal 4-qubit experiment. For a non-interacting Hamiltonian $H=Z_{0}+Z_{1}$, the evolution phase factors separably, yielding zero correlation between parameter qubits (Eq. 1). For an interacting Hamiltonian $H=Z_{0}Z_{1}$, the coupled phase creates entanglement, yielding $I(P_{0}:P_{1})=1.74$ bits (Eq. 2). This demonstrates that non-linearity requires interaction terms that couple parameter subspaces—a necessary condition for the trap-diffusion mechanism to operate.
@@ II.1 The Trap-Diffusion Mechanism
The core insight connecting my protocol to quantum walk speedup lies in the controlled-operation structure. In my W-gate protocol (see Methods), the circuit applies:
@@
This mechanism is mechanistically analogous to Grover’s amplitude amplification [grover1996fast], where the oracle marks “good” states (low energy) and the diffuser redistributes amplitude. My protocol realizes this as a continuous process: the controlled-mixer acts as the diffuser, and the energy-dependent sensing acts as a soft oracle that preferentially marks high-energy states for trapping.
Crucially, I observe that the work extraction correlates with Logarithmic Negativity ($R\approx 0.9$), a strict entanglement monotone. A purely classical feedback loop (zero entanglement) would yield zero work in this protocol, confirming the quantum nature of the mechanism.
To rigorously characterize the thermodynamic cost of this non-linearity, I implemented a controlled “Coherent Feedback” protocol (see Methods) designed to decouple the information gathering phase from the feedback actuation. Unlike standard variational updates where parameters are modified classically, my protocol utilizes a coherent conditional rotation ($CR_{X}$) acting directly on the system, triggered by the ancilla’s state.
Crucially, I fixed the feedback actuation strength (the “kick”) to a constant value and varied only the sensing duration $dt$. This isolates the information content as the sole variable driving the performance difference. I observe that the extracted work ($\Delta\langle H\rangle$) scales linearly with the mutual information $I(S:A)$ established between the system and the ancilla, yielding a strict constitutive Equation of State:
@@
Here, the proportionality constant $\eta$ represents the “Algorithmic Efficiency” or informational conductance of the ansatz-Hamiltonian pair. My data indicates a robust linear correlation ($R^{2}\approx 0.89$), confirming that the optimization rate is fundamentally limited by the information channel capacity of the probe, rather than the driving force of the optimizer.
Figure 1: The Thermodynamic Constitutive Law. Extracted Work ($\Delta\langle H\rangle$) vs. Mutual Information $I(S:A)$ for a 4-qubit transverse Ising system. The robust linear relationship ($R^{2}\approx 0.89$, slope $\eta\approx 0.11$ energy/bit) confirms the constitutive law $\Delta E\leq\eta I$. The sensing time $\tau$ was varied from 0 to 1.5 while the feedback strength was held constant at $\theta_{\text{gain}}=0.2$ rad (Table I), isolating information as the sole variable driving work extraction.
To determine the physical nature of these correlations, I measured the Logarithmic Negativity $E_{N}$, a strict monotone of quantum entanglement which is zero for all separable (classical) states. I find that the work extraction correlates directly with the generated entanglement:
@@
The persistence of this correlation ($R\approx 0.89$) and the strictly non-zero value of $E_{N}$ confirms that the mechanism relies specifically on quantum interference effects. This distinguishes the VQA optimizer from a classical Szilard engine, establishing that quantum entanglement is the essential thermodynamic resource driving the cooling process.
To verify that the protocol operates in the quantum regime, I computed the Landauer erasure cost. I observe a constant ratio $I(S:A)/S(A)=2.00$ across all sensing times (Fig. 2). This ratio of exactly 2 is the mathematical signature of pure bipartite entanglement, where $I(S:A)=2S(A)$ by definition. For any pure state $|\psi\rangle_{SA}$, the joint entropy $S(SA)=0$, yielding $I=S(S)+S(A)=2S(A)$.
Figure 2: Landauer Analysis of the Quantum Engine. Extracted work (blue) vs. Landauer erasure cost (red dashed) as a function of sensing time $\tau$. The constant ratio $I(S:A)/S(A)=2.00$ is the mathematical signature of pure bipartite entanglement ($I=2S$ by definition for pure states), confirming that the feedback mechanism operates via quantum correlations rather than classical correlations (which would yield $I\leq S$).
This confirms that the information channel operates via genuine quantum correlations rather than classical correlations (where $I\leq S$). The factor-of-2 ratio serves as a regime confirmation that the feedback mechanism utilizes entanglement. However, the absolute thermodynamic work extracted in my protocol does not exceed the Landauer erasure cost $W_{\text{cost}}=k_{B}T\ln(2)\cdot S(A)$ when the feedback coupling $\theta_{\text{gain}}$ is small. This is because my constitutive relation $W=\eta\cdot I(S:A)$ with $\eta\approx 0.11$ at $\theta_{\text{gain}}=0.2$ (Table I) yields less work per bit of information than the erasure cost of $\ln(2)\approx 0.69$ per bit. Net positive work extraction (where $W>W_{\text{cost}}$) would require stronger feedback coupling.
@@ III The Complexity-Dependent Efficiency Transition
To investigate the limits of this feedback mechanism, I analyzed the scaling of the algorithmic efficiency $\eta$ across two distinct topological classes of Hamiltonians as a function of system size $N$. I compared an “Ordered” system (Complete Graph $K_{n}$) characterized by a polynomial Dynamical Lie Algebra (DLA) dimension ($O(N^{3})$) [allcock2024dynamical] against a “Chaotic” system (Sherrington-Kirkpatrick Spin Glass) characterized by an exponential DLA dimension ($O(4^{N})$) [larocca2023diagnosing]. While the system sizes ($N\leq 8$) are limited by the exponential cost of classical simulation, the observed efficiency collapse in chaotic systems is consistent with the ancilla channel capacity becoming insufficient to resolve the scrambled gradient information.
The Lie algebraic classification of trainability was recently unified by Ragone et al. [ragone2024lie], who proved the exact variance formula $\text{Var}[\ell]=\mathcal{P}_{\mathfrak{g}}(\rho)\cdot\mathcal{P}_{\mathfrak{g}}(O)/\dim(\mathfrak{g})$, establishing that $\dim(\mathfrak{g})\in\Omega(b^{n})$ with $b>2$ implies a barren plateau. My efficiency metric $\eta$ provides a complementary thermodynamic diagnostic: while Ragone et al. predict trainability from gradient variance, I directly measure the information-to-work conversion capacity of the feedback loop. Both observables collapse when DLA dimension grows exponentially, suggesting a unified information-theoretic origin for barren plateaus.
I observe a sharp bifurcation in thermodynamic behavior (Fig. 3). For the Ordered system, I observe a regime of “Constructive Information Scaling,” where the efficiency $\eta$ increases with system size, indicating that the computational resistance to information flow decreases as the Hilbert space grows. In this regime, the growing algebraic structure provides additional pathways for the Demon to navigate the Hilbert space without exceeding its information channel capacity.
Figure 3: The Complexity-Dependent Efficiency Transition. Normalized algorithmic efficiency ($\eta/N^{2}$) vs. System Size NN for Ordered (Ferromagnet, blue) and Chaotic (Spin Glass, red) Hamiltonians. The Ordered phase maintains positive efficiency with non-monotonic finite-size oscillations [caleca2024giant], while the Chaotic phase undergoes efficiency collapse at moderate NN. Error bars represent $\pm 1\sigma$ over 5 random seeds. The critical system size $N_{c}$ depends on the sensing time $\tau$: longer sensing allows more information scrambling in chaotic systems, inducing earlier collapse.
@@
In stark contrast, the Chaotic system undergoes an “Efficiency Collapse.” As the system size increases, the efficiency drops precipitously, approaching zero at a $\tau$-dependent critical size $N_{c}$. At this point, the rate of operator spreading (scrambling) generates new independent gradient directions faster than the single-qubit ancilla (1 bit/cycle) can interrogate them. Notably, $N_{c}$ decreases with longer sensing times $\tau$, as the chaotic Hamiltonian has more time to scramble the gradient information before the ancilla can capture it. Beyond $N_{c}$, the information generation rate exceeds the channel capacity.
To quantify this transition, I introduce the Complexity Specific Heat, defined as the susceptibility of algorithmic efficiency to system size increase:
@@
This quantity is analogous to thermal specific heat $C=\partial\langle E\rangle/\partial T$, measuring the “response” of the optimization engine to changes in problem complexity. In the Ordered phase, $\chi_{\text{comp}}>0$ (efficiency increases with $N$), while in the Chaotic phase, $\chi_{\text{comp}}<0$ and diverges at the $\tau$-dependent critical size, signaling a thermodynamic instability.
This efficiency collapse provides a complementary diagnostic to the Quantum Fisher Information approach of Abbas et al. [abbas2021effective], who showed that the effective dimension $d_{\text{eff}}$ (derived from QFIM eigenvalues) correlates with model trainability. While their metric captures parameter redundancy, my efficiency $\eta$ directly measures the information-to-work conversion capacity of the feedback loop.
@@
For the Ordered phase, I utilized the Complete Graph Hamiltonian ($K_{n}$). Recent algebraic analysis by Allcock et al. [allcock2024dynamical] proves the exact dimension $\dim(\mathfrak{g}_{K_{n}})=\frac{1}{12}(n^{3}+6n^{2}+O(n))$, placing it firmly within the polynomially tractable regime.